# MeshGraphNet-Transformer: Scalable Mesh-based Learned Simulation for Solid Mechanics

### v1

@@ 4.3 Computational Cost
All experiments were executed on a single NVIDIA RTX 4090 GPU with 24 GB of memory. Inference time consumption is described in Table 6, where $t_{\text{model}}$ represents the time consumption of one forward pass and $t_{\text{full}}$ includes the preprocessing for each time step. Highlight that, for the pi-beam dataset, MGN suffers from under-reaching. Training MGN-T required approximately 4 hours on the pi-beam dataset and 8 hours on the DeformingPlate dataset, highlighting the computational efficiency of the proposed approach.
@@ 5 Discussion
On the deforming plate dataset, MGN-T achieves competitive results compared to state-of-the-art approaches, demonstrating low error propagation in both single-step and full-rollout predictions. Compared to other approaches, MGN-T requires a fraction (roughly 1/4) of the parameters, making it a much lighter model , being its training and inference fast and efficient. These results confirm that the proposed method is a promising off-the-shelf approach for simulating solid mechanics phenomena, as it does not require form pre-processing, coarsening, or segmentation, unlike other hierarchical graph-based approaches.

### v3

@@ 4.3 Computational Cost
All experiments were executed on a single NVIDIA RTX 4090 GPU (24 Gb) of memory. Inference time consumption is described in Table 6, where $t_{\text{model}}$ represents the time consumption of one forward pass and $t_{\text{full}}$ includes the preprocessing for each time step. Highlight that, for the pi-beam dataset, MGN suffers from under-reaching. Training MGN-T required approximately 4 hours on the Pi-beam dataset and 8 hours on the Deforming plate dataset, highlighting the computational efficiency of the proposed approach.
@@ 5 Discussion
On the deforming plate dataset, MGN-T achieves competitive results compared to state-of-the-art approaches, demonstrating low error propagation in both single-step and full-rollout predictions. Compared to other approaches, MGN-T requires a fraction (roughly 1/4) of the parameters, making it a much lighter model, being its training and inference fast and efficient. These results confirm that the proposed method is a promising off-the-shelf approach for simulating solid mechanics phenomena, as it does not require any pre-processing, coarsening, or segmentation, unlike other hierarchical graph-based approaches.